### v1

@@ 3 Several propositions
can be distinguished by subwords of length $23m+$ 2, $33m+$ 3, $53m+$ 5 and $33m+$ 3, respectively, and not smaller than that.
@@
Consider the words $u=0^{m-1}10^{m}10^{m+1}10^{m+1}1$ and $v=0^{m}10^{m-1}10^{m+1}10^{m+1}1$. Let $w$ be a subword of exactly one of the words $u$ and $v$. We need to prove that $|w|\geqslant 3m+5$. Without loss of generality we can assume that $w$ is a subword of the word $u$ and is not a subword of $v$ (the case when $w$ is a subword of $v$ and not a subword of $u$ is symmetric, since $u$ and $v$ are mirror images of each other). Since the word $10^{m-1}10^{m-1}10^{m+1}10^{m+1}$ 1 is a subword of $v$, then $w$ must include the second block $0^{m}$ of $u$. Indeed, if $w$ does not include the second block $0^{m}$, then $w$ is a subword of $10^{m-1}10^{m-1}10^{m+1}10^{m+1}$ 1, and hence it is a subword of $v$. Similarly, since the words $10^{m-1}10^{m}10^{m-1}10^{m+1}$ 1 and $10^{m-1}10^{m}10^{m+1}10^{m}$ 1 are subwords of $v$, then $w$ must contain the fourth block $0^{m+1}$ of $u$ and must contain at least $m$ zeros from the third block $0^{m+1}$ of $u$. Besides that, $w$ must have at least three 1’s, because each subword of $u$ which has at most two occurrences of 1’s is a subword of $v$. So, $|w|\geqslant(m+(m+1)+m+3)+1=3m+5$. To finish the proof, it remains to notice that the word $110^{2m+1}10^{m+1}$ 1 of length $53m+$ 5 is a subword of the word $u$ and is not a subword of $v$.
For other cases we provide words which are subwords of $u$ and is not subwords of $v$. For $n=4m+3$ one can take the subword $110^{m-1}10^{m-1}10^{m}$ 1 of length $23m+$ 2, for $n=4m+4$ the subword $110^{m-1}10^{m-1}10^{m+1}$ 1 of length $33m+$ 3, $n=4m+6$ the subword $110^{m-1}10^{m-1}10^{m+1}$ 1 of length $33m+$ 3.
The following proposition and corollary are the key tools for the proof of the theorem for the case of non-special and nonperiodic words (Lemma 4.2).
@@
Let $u=01^{\alpha_{1}}01^{\alpha_{2}}\cdots 01^{\alpha_{l}}$ be a cyclic word which is not special and not periodic, $l\geqslant 2$, and $\alpha_{i}\geqslant 1$ for each $1\leqslant i\leqslant l$. Then there exists a unique subword $u_{fix}$ of $u$ which contains $n_{1,u}=n_{1}$ occurrences of 1 (all 1’s from $u$) and at most $\frac{l+2}{2}$ occurrences of 0.
@@
Let $k$ be the minimal length of blocks of 1’s: $k=\min_{1\leqslant i\leqslant l}\alpha_{i}$, and let $s$ denote the maximal integer such that $(01^{k})^{s}0$ is a contiguous subword of $u$. So, $u=(01^{k})^{s}01^{\alpha_{1}}\cdots 01^{\alpha_{r}}$, where $s+r=l$, $\alpha_{i}\geqslant k$ for each $i$. Since $s$ is the maximal possible, we have in particular that $\alpha_{1},\alpha_{r}>k$. Since $u$ is not special, then $s<l$. Consider two cases.
1. $s\leqslant\frac{l}{2}$.
@@
Since $s$ is the maximal possible, then for any two subwords equal to $u^{\prime}$, there is no occurrence of 00 that is included in both subwords. Words $u$ and $u^{\prime}$ have $l$ and $s+1$ occurrences of 0, respectively. So, there are at most $\frac{l}{s+1}$ ways to take a subword of $u$ which is equal to $u^{\prime}$.
Consider a subword $w$ in $u$ which equal to $u^{\prime}$. Since $u$ is not periodic, then for every unit turn of $u$ which translates $u^{\prime}$ to $w$ we can find a zero in $u$ which is translated to an empty place. Consider a word $w^{\prime}$ obtained by adding all these 0’s to $w$. Since there are at most $\frac{l}{s+1}$ ways to take a subword of $u$ which is equal to $u^{\prime}$, then $w^{\prime}$ has $n_{1}$ occurrences of 1 and at most $f(s)=s+1+(\frac{l}{s+1}-1)=s+\frac{k}{s+1}$ occurrences of 0, and $w^{\prime}$ is unique. For $1\leqslant s\leqslant\frac{l-2}{2}$ we have $w^{\prime}_{0}\leqslant s+\frac{l}{s+1}\leqslant\frac{l+2}{2}$. For $\frac{l-2}{2}<s\leqslant\frac{l}{2}$, the subword $u^{\prime}$ is unique and $u^{\prime}_{0}=s+1\leqslant\frac{l+2}{2}$. In both cases we find a unique subword with $n_{1}$ occurrences of 1 and at most $\frac{l+2}{2}$ occurrences of 0.
2. $s>\frac{l}{2}$.
Since $u=(01^{k})^{s}01^{\alpha_{1}}\ldots 01^{\alpha_{r}}$ is not special, then three cases are possible:
  • $k\nmid\alpha_{i}$ for some $i$ (case 2.0)
  • $u=(01^{k})^{s}01^{k\beta_{1}}01^{k\beta_{2}}\cdots 01^{k\beta_{r}}$ and $\beta_{1}\geqslant 3$ (or $\beta_{r}\geqslant 3)$ (case 2.1)
  • $u=(01^{k})^{s}01^{k\beta_{1}}01^{k\beta_{2}}\cdots 01^{k\beta_{r}}$, $\beta_{1}=\beta_{r}=2$ and $\beta_{i}\geqslant 2$ for some $1<i<r$ (case 2.2)
We now consider these three subcases.
2.0 $k\nmid\alpha_{i}$.
Let $1^{t}$ be the smallest block of 1’s such that $k\nmid t$. Then $t$ cannot be equal to a sum of two or more integers from $\{k,\alpha_{1},\ldots,\alpha_{r}\}$. Let $s^{\prime}$ be the maximal number of consecutive blocks $1^{t}$. Since $s>\frac{l+2}{2}$, then $s^{\prime}<\frac{l-2}{2}$ Consider the word $u^{\prime}=(01^{t})^{s^{\prime}}01^{n_{1}-ts^{\prime}}$. There are at most $\frac{l}{s^{\prime}+1}$ ways to take a subword of $u$ which is equal to $u^{\prime}$. So, with an argument from Case 1 we prove that we can find a unique subword with $n_{1}$ occurrences of 1 and at most $\frac{l+2}{2}$ occurrences of 0.
@@
2.1 $u=(01^{k})^{s}01^{k\beta_{1}}01^{k\beta_{2}}\cdots 01^{k\beta_{r}}$ and $\beta_{1}\geqslant 3$ (or $\beta_{r}\geqslant 3)$
Without loss of generality we assume that $\beta_{1}\geqslant 3$. Consider the word
$u^{\prime}=(01^{k})^{r}(1^{k}01^{k})^{\frac{s-r}{2}}01^{k\beta_{1}}1^{k\beta_{2}}\cdots 1^{k\beta_{r}}\quad\text{if}\quad 2\mid(s-r)$
$u^{\prime}=(01^{k})^{r}(1^{k}01^{k})^{\frac{s-r-1}{2}}1^{k}01^{k\beta_{1}}1^{k\beta_{2}}\cdots 1^{k\beta_{r}}\quad\text{if}\quad 2\nmid(s-r).$
[…]
In both cases $u^{\prime}$ has $n_{1}$ occurrences of 1 and at most $r+\frac{s-r}{2}+1=\frac{s+r+2}{2}\leqslant\frac{l+2}{2}$ occurrences of 0.
Now we prove that $u^{\prime}$ is unique for the case $2\mid(s-r)$ (the second case is similar). Assume the converse: suppose that there exists a subword $u^{\prime\prime}=u^{\prime}$ and a unit turn $\sigma$ of $u$ which translates $u^{\prime}$ to $u^{\prime\prime}$. We now index 00’s in $u$: $u=0_{1}1^{k}0_{2}1^{k}\cdots 0_{s}1^{k}0_{s+1}1^{k\beta_{1}}0_{s+2}1^{k\beta_{2}}\cdots 0_{s+r}1^{k\beta_{r}}$. If $\sigma(0_{1})=0_{2}$, then $\sigma(0_{s+1})=\varnothing$, a contradiction. If $\sigma(0_{1})=0_{3}$ or $\sigma(0_{1})=0_{4}$, then $\sigma(0_{s})=\varnothing$, a contradiction. If $\sigma(0_{1})=0_{5}$ or $\sigma(0_{1})=0_{6}$, then $\sigma(0_{s-2})=\varnothing$, a contradiction. Continuing this line of reasoning, we get that $\sigma(0_{1})\neq 0_{2},0_{3},\ldots,0_{s+1}$. Similarly, $\sigma(0_{1})\neq 0_{s+2},0_{s+3},\ldots,0_{s+r}$, since in this case there exists $11\leqslant i\leqslant r-$ 1, such that $\sigma(0_{i})=0_{s+r}$ and $\sigma(0_{i+1})=0_{1}$. Recall that the number of 1’s between $0_{i}$ and $0_{i+1}$ is equal to $k$, and $\beta_{r}>k$ is the number of 1’s between $0_{s+r}$ and $0_{1}$. A contradiction, since both $u^{\prime}$ and $u^{\prime\prime}$ have $n_{1}$ occurrences of 1 and $0_{1},0_{2},\ldots,0_{r}$. So, $\sigma$ is the identical turn, and hence $u^{\prime}$ is a unique subword.
2.2 $u=(01^{k})^{s}01^{k\beta_{1}}01^{k\beta_{2}}\cdots 01^{k\beta_{r}}$, $\beta_{1}=\beta_{r}=2$ and $\beta_{i}\geqslant 2$ for some $1<i<r$.
@@
where $a,b\geqslant 1$, $\beta_{b+1},\beta_{j},\beta_{j+a+1}>1$, and $a$ is the maximal possible. So, there are $b-1$ consecutive blocks $1^{k}$ after the first block $1^{2k}$, and $a-1$ is the maximal number of consecutive blocks $1^{k}$ in the remaining part. Without of loss generality we can assume that $a\geqslant b$ (indeed, otherwise we can take a mirror image of $u$: $(01^{k})^{s}01^{k\beta_{r}}01^{k\beta_{r-1}}\cdots 01^{k\beta_{1}}$). Also we can suppose that parts $(01^{k})^{b-1}$ and $(01^{k})^{a-1}$ do not coincide, since there exists $\beta_{i}\geqslant 2$. Let $t$ be an integer such that $s=(a+1)+t(2b+2)+x$, where $x<2b+2$.
@@
$u^{\prime}=(01^{k})^{a+1}((1^{k})^{b+1}(01^{k})^{b+1})^{t}(01^{k})^{x}01^{k\beta_{1}}1^{k\beta_{2}}\cdots 1^{k\beta_{r}},\quad\text{if}\quad x\leqslant b$
$u^{\prime}=(01^{k})^{a+1}((1^{k})^{b+1}(01^{k})^{b+1})^{t}(1^{k})^{x-b}(01^{k})^{b}01^{k\beta_{1}}1^{k\beta_{2}}\cdots 1^{k\beta_{r}},\quad\text{if}\quad b<x<2b+2.$
[…]
@@
$=2(a+1)+2t(b+1)+(b+x)\geqslant 2(a+1+t(b+1)+\min(b,x)+1)-2.$
We now prove that $u^{\prime}$ is unique in the case $2\mid(s-r)$ (in the second case the proof is similar). Assume the converse, i.e. suppose that there exists a subword $u^{\prime\prime}=u^{\prime}$ and a unit turn of $u$ which translates $u^{\prime}$ to $u^{\prime\prime}$. We index 00’s and $1$’s in $u$ as follows:
@@
If $\sigma(0_{1})=0_{s-(a-2)-t^{\prime}(2b+2)-x^{\prime}}$ and $s-(a-2)-t^{\prime}(2b+2)-x^{\prime}>0$ for $11\leqslant x^{\prime}\leqslant b+$ 1, then $\sigma(0_{a+1+t^{\prime}(2b+2)+(b+1)+x^{\prime}})=\varnothing$ and $0_{a+1+t^{\prime}(2b+2)+(b+1)+x^{\prime}}\in u^{\prime}$. A contradiction.
@@
Let $u=0^{x}1^{\alpha_{1}}0^{x}1^{\alpha_{2}}\cdots 0^{x}1^{\alpha_{l}}$ be a cyclic word which is not special and not periodic, $l\geqslant 2$, and $\alpha_{i}\geqslant 1$ for each $1\leqslant i\leqslant l$. Then there exists a unique subword $u_{fix}$ of $u$ which contains $n_{1,u}=n_{1}$ occurrences of 1 (all 1’s from $u$) and at most $\frac{l+2}{2}$ blocks $0^{x}$.
Proof 3.6.
The previous proposition implies that the word $01^{\alpha_{1}}01^{\alpha_{2}}\cdots 01^{\alpha_{l}}$ has a unique subword $01^{\beta_{1}}01^{\beta_{2}}\ldots 01^{\beta_{r}}$, which has $n_{1}$ occurrences of 1 and at most $\frac{l+2}{2}$ occurrences of 0. Then the word $0^{x}1^{\beta_{1}}0^{x}1^{\beta_{2}}\cdots 0^{x}1^{\beta_{r}}$ is a unique subword of $u$.
@@
Proof 3.8.
Without loss of generality assume that the number of 1’s in $u$ is less or equal to the number of 0’s (we can check which letter is more frequent by checking if the word $10^{\lfloor\frac{n}{2}\rfloor}+$ 1 is a factor). Then $S$ contains the word $1^{n_{1}}$ and does not contain the word $1^{n_{1}+1}$. So we can recover $n_{1}$, and $n_{0}=n-n_{1}$.
It is clear that $l\leqslant n_{1}\leqslant\frac{n}{2}$. If $l\leqslant\frac{3}{8}n,$ then $S$ contains the word $(01)^{l+1}$. In this case we can find the word $(01)^{k}$ in $S$ with maximal $k$ and recover $l$.
If $l>\frac{3}{8}n$, then we can consider the subword $u^{\prime}=0^{a_{1}}1^{b_{1}}\cdots 0^{a_{k}}1^{b_{k}}$ of word $u$, which has $n_{0}$ occurrences of 0 and each block of 1’s which contains at least two occurrences of 1. Let $l_{1}$ be the number of blocks 1’s which length is 1. Then $l_{1}+k=l\geqslant\frac{3}{8}n$ and $l_{1}+2k\leqslant n_{1}\leqslant\frac{n}{2}$. So $\frac{3}{4}n\leqslant 2l_{1}+2k\leqslant l_{1}+n_{1}\leqslant l_{1}+\frac{n}{2}$. It means that $l_{1}\geqslant\frac{n}{4}$ and $|u^{\prime}|=n-l_{1}\leqslant\frac{3}{4}n$. So we can find in $S$ the word with $n_{1}$ occurrences of 1: $0^{a_{1}}1^{b_{1}}\cdots 0^{a_{k}}1^{b_{k}}$, where $b_{i}\geqslant 2$ for all $1\leqslant i\leqslant k$ and $k,b_{i}$ — maximal possible. It is word $u^{\prime}$. So we can recover $l$: $l=k+l_{1}=k+(n-|v|)$.
Corollary 3.9.
Let $u$ and $v$ be two cyclic words of length $n$. If $n_{0,u}\neq n_{0,v}$ or $n_{1,u}\neq n_{1,v}$ or $l_{u}\neq l_{v}$, then there is a word of length at most $\frac{3}{4}n+4$ which is a subword of only one of words $u$ and $v$.
Proof 3.10.
Follows from Proposition 3.7.
In the proof of the main result, for finding a subword of only one of the words $u$ and $v$, we will often use a technique described in the following proposition:
Proposition 3.11.
Let $u\neq v$ be two cyclic words such that $n_{0,u}=n_{0,v}$, $n_{1,u}=n_{1,v}=n_{1}$. Suppose that $u^{\prime}$ is a subword of $u$ which contains $n_{1}$ occurrences of 1 and that there is only identical unit turn which translates $u^{\prime}$ to the same subword. Suppose in addition that $|u^{\prime}|+y+1\leqslant\frac{3}{4}n+4$, where $y$ is the length of the longest block in $u$ which is shorter than $0^{x}$. Then there is a subword of only one of the words $u$ and $v$ of length at most $\frac{3}{4}n+4$.

### v2

@@ 3 Preliminary observations and auxiliary statements
can be distinguished by subwords of length $3m+2$, $3m+3$, $3m+5$ and $3m+3$, respectively, and not smaller than that.
@@
Consider the words $u=0^{m-1}10^{m}10^{m+1}10^{m+1}1$ and $v=0^{m}10^{m-1}10^{m+1}10^{m+1}1$. Let $w$ be a distinguishing subword for $u$ and $v$. We need to prove that $|w|\geqslant 3m+5$. Without loss of generality we can assume that $w$ is a subword of the word $u$ and is not a subword of $v$ (the case when $w$ is a subword of $v$ and not a subword of $u$ is symmetric, since $u$ and $v$ are mirror images of each other as cyclic words). Since the word $0^{m-1}10^{m-1}10^{m+1}10^{m+1}1$ is a subword of $v$, $w$ must include the second block $0^{m}$ of $u$. Indeed, if $w$ does not include the second block $0^{m}$, then $w$ is a subword of $0^{m-1}10^{m-1}10^{m+1}10^{m+1}1$, and hence it is a subword of $v$. Similarly, since the words $0^{m-1}10^{m}10^{m+1}10^{m}1$ and $0^{m-1}10^{m}10^{m-1}10^{m+1}1$ are subwords of $v$, then $w$ must contain the fourth block $0^{m+1}$ of $u$ and must contain at least $m$ zeros from the third block $0^{m+1}$ of $u$. Besides that, $w$ must have at least three 1’s, because each subword of $u$ which has at most two occurrences of 1 is a subword of $v$. So, $|w|\geqslant(m+(m+1)+m+3)+1=3m+5$. To finish the proof, it remains to notice that the word $10^{2m+1}10^{m+1}1$ of length $3m+5$ is a subword of the word $u$ and is not a subword of $v$.
For other cases we provide words which are subwords of $u$ and are not subwords of $v$. For $n=4m+3$ one can take the subword $10^{m-1}10^{m-1}10^{m}1$ of length $3m+2$, for $n=4m+4$ the subword $10^{m-1}10^{m-1}10^{m+1}1$ of length $3m+3$, $n=4m+6$ the subword $10^{m-1}10^{m-1}10^{m+1}1$ of length $3m+3$. ∎
@@
Without loss of generality assume that the number of 1’s in $u$ is less than or equal to the number of 0’s (we can check which letter is more frequent by checking if the word $0^{\lfloor\frac{n}{2}\rfloor+1}$ is a subword). Then $S$ contains the word $1^{n_{1}}$ and does not contain the word $1^{n_{1}+1}$. So we can recover $n_{1}$, and $n_{0}=n-n_{1}$.
It is clear that $l\leqslant n_{1}\leqslant\frac{n}{2}$. If $l\leqslant\frac{3}{8}n,$ then $S$ contains the word $(01)^{l}$. In this case we can find the word $(01)^{k}$ in $S$ with maximal $k$ and recover $l$.
If $l>\frac{3}{8}n$, then we can consider the subword $u^{\prime}=0^{a_{1}}1^{b_{1}}\cdots 0^{a_{k}}1^{b_{k}}$ of the word $u$, which has $n_{0}$ occurrences of 0 and each block of 1’s containing at least two occurrences of 1. Let $l_{1}$ be the number of blocks 1 which have length 1. Then $l_{1}+k=l\geqslant\frac{3}{8}n$ and $l_{1}+2k\leqslant n_{1}\leqslant\frac{n}{2}$. So, $\frac{3}{4}n\leqslant 2l_{1}+2k\leqslant l_{1}+n_{1}\leqslant l_{1}+\frac{n}{2}$. It means that $l_{1}\geqslant\frac{n}{4}$, and hence $|u^{\prime}|=n-l_{1}\leqslant\frac{3}{4}n$. Thus, $u^{\prime}$ is in $S$. Moreover, we can detect it in $S$ as a word of the form $0^{a_{1}}1^{b_{1}}\cdots 0^{a_{k}}1^{b_{k}}$, with $n_{0}$ occurrences of 0, $b_{i}\geqslant 2$ for all $1\leqslant i\leqslant k$, and with $k$ and $b_{i}$ being maximal. So, we can recover $l$ from $u^{\prime}$: $l=k+l_{1}=k+(n-|u^{\prime}|)$. ∎
Corollary 3.5.
Let $u$ and $v$ be two cyclic words of length $n$. If $n_{0,u}\neq n_{0,v}$ or $n_{1,u}\neq n_{1,v}$ or $l_{u}\neq l_{v}$, then there exists a distinguishing subword of length at most $\frac{3}{4}n+4$ for $u$ and $v$.
Proof 3.6.
Follows from Proposition 3.3. ∎
The following proposition and corollary are key tools for the proof of the main theorem for the case of neither special nor periodic words (we use it in Lemma 4.2).
Proposition 3.7.
Let $u=01^{\alpha_{1}}01^{\alpha_{2}}\cdots 01^{\alpha_{l}}$ be a cyclic word which is neither special nor periodic, with $l\geqslant 2$, and $\alpha_{i}\geqslant 1$ for each $1\leqslant i\leqslant l$. Then there exists a unioccurrent subword $u_{\operatorname{fix}}$ of $u$ which contains $n_{1,u}=n_{1}$ occurrences of 1 (all 1’s from $u$) and at most $\frac{l+2}{2}$ occurrences of 0.
Proof 3.8.
Let $k$ be the minimal length of blocks of 1’s: $k=\min_{1\leqslant i\leqslant l}\alpha_{i}$, and let $s$ denote the maximal integer such that $(01^{k})^{s}0$ is a factor of $u$. So, $u=(01^{k})^{s}01^{\alpha_{1}}\cdots 01^{\alpha_{r}}$, where $s+r=l$, $\alpha_{i}\geqslant k$ for each $i$. Since $s$ is chosen maximal, we have in particular that $\alpha_{1},\alpha_{r}>k$. Since $u$ is not special, we have $s<l$. There are two cases to consider:
Case 1. $s\leqslant\frac{l}{2}$.
@@
Since $s$ is chosen maximal, for any two occurrences of subwords equal to $u^{\prime}$, there is no occurrence of 0 that is included in both subwords. The words $u$ and $u^{\prime}$ have $l$ and $s+1$ occurrences of 0, respectively. So, there are at most $\frac{l}{s+1}$ ways to take an occurrence of a subword of $u$ which is equal to $u^{\prime}$.
Consider an occurrence of a subword $w$ in $u$ which is equal to $u^{\prime}$. Since $u$ is not periodic, for every turn of $u$ which translates $u^{\prime}$ to $w$ we can find a zero in $u$ which is translated to an empty place. Consider a word $w^{\prime}$ obtained by adding all these 0’s to $w$. Since there are at most $\frac{l}{s+1}$ ways to take an occurrence of a subword of $u$ which is equal to $u^{\prime}$, $w^{\prime}$ has $n_{1}$ occurrences of 1 and at most $s+1+(\frac{l}{s+1}-1)=s+\frac{l}{s+1}$ occurrences of 0, and $w^{\prime}$ is unioccurrent. For $1\leqslant s\leqslant\frac{l-2}{2}$ we have $w^{\prime}_{0}\leqslant s+\frac{l}{s+1}\leqslant\frac{l+2}{2}$. For $\frac{l-2}{2}<s\leqslant\frac{l}{2}$, the subword $u^{\prime}$ is unioccurrent and $u^{\prime}_{0}=s+1\leqslant\frac{l+2}{2}$. In both cases we find a unioccurrent subword with $n_{1}$ occurrences of 1 and at most $\frac{l+2}{2}$ occurrences of 0.
Case 2. $s>\frac{l}{2}$.
We spit this case into three subcases as follows. Case 2.1 corresponds to $k\nmid\alpha_{i}$ for some $i$. If $k\mid\alpha_{i}$, we distinguish between two cases: either $u=(01^{k})^{s}01^{k\beta_{1}}01^{k\beta_{2}}\cdots 01^{k\beta_{r}}$ and $\beta_{1}\geqslant 3$ (or $\beta_{r}\geqslant 3)$ (Case 2.2), or $u=(01^{k})^{s}01^{k\beta_{1}}01^{k\beta_{2}}\cdots 01^{k\beta_{r}}$, $\beta_{1}=\beta_{r}=2$ (Case 2.3); in the latter case we have $\beta_{i}\geqslant 2$ for some $1<i<r$ since $u$ is not special. We now consider these three subcases.
Case 2.1. $k\nmid\alpha_{i}$.
Let $1^{t}$ be the smallest block of 1’s such that $k\nmid t$. Then $t$ cannot be equal to the sum of two or more integers from $\{k,\alpha_{1},\ldots,\alpha_{r}\}$. Let $s^{\prime}$ be the maximal number of consecutive blocks $1^{t}$. Since $s>\frac{l}{2}$, we have $s^{\prime}<\frac{l}{2}$. Consider the word $u^{\prime}=(01^{t})^{s^{\prime}}01^{n_{1}-ts^{\prime}}$. There are at most $\frac{l}{s^{\prime}+1}$ ways to take an occurrence of a subword of $u$ which is equal to $u^{\prime}$. So, with an argument similar to Case 1 we prove that we can find a unioccurrent subword with $n_{1}$ occurrences of 1 and at most $\frac{l+2}{2}$ occurrences of 0.
@@
Case 2.2. $u=(01^{k})^{s}01^{k\beta_{1}}01^{k\beta_{2}}\cdots 01^{k\beta_{r}}$ and $\beta_{1}\geqslant 3$ (or $\beta_{r}\geqslant 3)$
Without loss of generality we assume that $\beta_{1}\geqslant 3$. Recall that in Case 2 we have $s>r$, since $s>l/2$. Consider the word
$u^{\prime}=\begin{cases}(01^{k})^{r}(1^{k}01^{k})^{\frac{s-r}{2}}01^{k\beta_{1}}1^{k\beta_{2}}\cdots 1^{k\beta_{r}},&\text{if }2\mid(s-r),\\ (01^{k})^{r}(1^{k}01^{k})^{\frac{s-r-1}{2}}1^{k}01^{k\beta_{1}}1^{k\beta_{2}}\cdots 1^{k\beta_{r}},&\text{if }2\nmid(s-r).\end{cases}$
In both cases $u^{\prime}$ has $n_{1}$ occurrences of 1 and at most $r+\frac{s-r}{2}+1=\frac{s+r+2}{2}=\frac{l+2}{2}$ occurrences of 0. For example, if $u=\underline{0}1\underline{0}1\underline{0}1\,01\underline{0}1\,01\underline{0}1\,\underline{0}\,1^{3}0101^{2}$ ($r=3$, $s=7$, $k=1$), then $u^{\prime}$ contains all $1$’s and underlined 0’s: $u^{\prime}=\underline{0}1\underline{0}1\underline{0}1\,1\underline{0}1\,1\underline{0}1\,\underline{0}\,1^{3}11^{2}$. As another example, take $u=\underline{0}1\underline{0}1\underline{0}1\,01\underline{0}1\,01\underline{0}1\,01\underline{0}\,1^{3}0101^{2}$ ($r=3$, $s=8$, $k=1$); then $u^{\prime}$ contains all $1$’s and underlined 0’s: $u^{\prime}=\underline{0}1\underline{0}1\underline{0}1\,1\underline{0}1\,1\underline{0}1\,1\underline{0}\,1^{3}11^{2}$.
Now we prove that $u^{\prime}$ is unioccurrent in the case $2\mid(s-r)$ (the case $2\nmid(s-r)$ is similar). Assume the converse: suppose that there exists another occurrence $u^{\prime\prime}$ of the subword $u^{\prime}$, i.e. $u^{\prime\prime}=u^{\prime}$ and a turn $\sigma$ of $u$ which translates $u^{\prime}$ to $u^{\prime\prime}$. We now index 0’s in $u$: $u=0_{1}1^{k}0_{2}1^{k}\cdots 0_{s}1^{k}0_{s+1}1^{k\beta_{1}}0_{s+2}1^{k\beta_{2}}\cdots 0_{s+r}1^{k\beta_{r}}$. If $\sigma(0_{1})=0_{2}$, then $\sigma(0_{s+1})=\varnothing$, a contradiction. If $\sigma(0_{1})=0_{3}$ or $\sigma(0_{1})=0_{4}$, then $\sigma(0_{s})=\varnothing$, a contradiction. If $\sigma(0_{1})=0_{5}$ or $\sigma(0_{1})=0_{6}$, then $\sigma(0_{s-2})=\varnothing$, a contradiction. Continuing this line of reasoning, we get that $\sigma(0_{1})\neq 0_{2},0_{3},\ldots,0_{s+1}$. Similarly, $\sigma(0_{1})\neq 0_{s+2},0_{s+3},\ldots,0_{s+r}$, since in this case there exists $1\leqslant i\leqslant r-1$, such that $\sigma(0_{i})=0_{s+r}$ and $\sigma(0_{i+1})=0_{1}$. Recall that the number of 1’s between $0_{i}$ and $0_{i+1}$ is equal to $k$, and $\beta_{r}>k$ is the number of 1’s between $0_{s+r}$ and $0_{1}$. We reach a contradiction, since both $u^{\prime}$ and $u^{\prime\prime}$ have $n_{1}$ occurrences of 1 and $0_{1},0_{2},\ldots,0_{r}$. So, $\sigma$ is the identical turn, and hence $u^{\prime}$ is a unioccurrent subword.
Case 2.3. $u=(01^{k})^{s}01^{k\beta_{1}}01^{k\beta_{2}}\cdots 01^{k\beta_{r}}$, $\beta_{1}=\beta_{r}=2$ and $\beta_{i}\geqslant 2$ for some $1<i<r$.
@@
where $a,b\geqslant 1$, $\beta_{b+1},\beta_{j},\beta_{j+a+1}>1$, and $a$ is chosen maximal. So, there are b1b-1 consecutive blocks $1^{k}$ after the first block $1^{2k}$, and a1a-1 is the maximal number of consecutive blocks $1^{k}$ in the remaining part. Without loss of generality we can assume that $a\geqslant b$ (indeed, otherwise we can take a mirror image of $u$: $(01^{k})^{s}01^{k\beta_{r}}01^{k\beta_{r-1}}\cdots 01^{k\beta_{1}}$). We can also suppose that the parts $(01^{k})^{b-1}$ and $(01^{k})^{a-1}$ do not coincide, since there exists $\beta_{i}\geqslant 2$. Let $t$ be an integer such that $s=(a+1)+t(2b+2)+x$, where $x<2b+2$.
@@
$u^{\prime}=\begin{cases}(01^{k})^{a+1}((1^{k})^{b+1}(01^{k})^{b+1})^{t}(01^{k})^{x}01^{k\beta_{1}}1^{k\beta_{2}}\cdots 1^{k\beta_{r}},&\text{ if }x\leqslant b,\\ (01^{k})^{a+1}((1^{k})^{b+1}(01^{k})^{b+1})^{t}(1^{k})^{x-b}(01^{k})^{b}01^{k\beta_{1}}1^{k\beta_{2}}\cdots 1^{k\beta_{r}},&\text{ if }b<x<2b+2.\end{cases}$
For example, if $u=010101010101010\,1^{2}01^{2}0101^{2}$ ($r=4$, $a=2$, $b=1$, $k=1$, $s=7=3+1\cdot 4+0$), then $u^{\prime}$ contains all 1’s from $u$ and underlined 0’s: $u=\underline{0}1\underline{0}1\underline{0}1\,0101\underline{0}1\underline{0}1\,\underline{0}\,1^{2}01^{2}0101^{2}$.
The word $u^{\prime}$ has $n_{1}$ occurrences of 1 and $a+1+t(b+1)+\min(b,x)+1\leqslant\frac{l+2}{2}$ occurrences of 0, since
@@
$2(a+1)+2t(b+1)+(b+x)\geqslant 2(a+1+t(b+1)+\min(b,x)+1)-2.$
We now prove that $u^{\prime}$ is unioccurrent in the case $2\mid(s-r)$ (in the case $2\nmid(s-r)$ the proof is similar). Assume the converse, i.e. suppose that there exists a subword $u^{\prime\prime}$ such that $u^{\prime\prime}=u^{\prime}$ and a turn of $u$ which translates $u^{\prime}$ to $u^{\prime\prime}$. We index 0’s and 1’s in $u$ as follows:
@@
If $\sigma(0_{1})=0_{s-(a-2)-t^{\prime}(2b+2)-x^{\prime}}$ and $s-(a-2)-t^{\prime}(2b+2)-x^{\prime}>0$ for $1\leqslant x^{\prime}\leqslant b+1$, then $\sigma(0_{a+1+t^{\prime}(2b+2)+(b+1)+x^{\prime}})=\varnothing$ and $0_{a+1+t^{\prime}(2b+2)+(b+1)+x^{\prime}}\in u^{\prime}$. A contradiction.
@@
Let $u=0^{x}1^{\alpha_{1}}0^{x}1^{\alpha_{2}}\cdots 0^{x}1^{\alpha_{l}}$ be a cyclic word which is neither special nor periodic, with $l\geqslant 2$, and $\alpha_{i}\geqslant 1$ for each $1\leqslant i\leqslant l$. Then there exists a unioccurrent subword $u_{\operatorname{fix}}$ of $u$ which contains $n_{1}$ occurrences of 1 (all 1’s from $u$) and at most $\frac{l+2}{2}$ blocks $0^{x}$ (and no other zeros).
Proof 3.10.
The previous proposition implies that the word $01^{\alpha_{1}}01^{\alpha_{2}}\cdots 01^{\alpha_{l}}$ has a unioccurrent subword $01^{\beta_{1}}01^{\beta_{2}}\ldots 01^{\beta_{r}}$, which has $n_{1}$ occurrences of 1 and at most $\frac{l+2}{2}$ occurrences of 0. Then the word $0^{x}1^{\beta_{1}}0^{x}1^{\beta_{2}}\cdots 0^{x}1^{\beta_{r}}$ is a unioccurrent subword of $u$. ∎
In the proof of the main result, for finding a distinguishing subword for the words $u$ and $v$, we often use a technique described in the following proposition:
@@
Let $u\neq v$ be two cyclic words such that $n_{0,u}=n_{0,v}$, $n_{1,u}=n_{1,v}=n_{1}$, ${x_{u}}={x_{v}}=x$. Suppose that $u^{\prime}$ is either a unioccurrent subword of $u$ or a subword of $u$ which contains $n_{1}$ occurrences of 1 and such that the only turn translating $u^{\prime}$ to an occurrence of an equal subword is the identity map. Suppose also that $|u^{\prime}|+y+1\leqslant\frac{3}{4}n+4$, where $y$ is the length of the longest block in $u$ which is shorter than $0^{x}$. Then there exists a distinguishing subword of length at most $\frac{3}{4}n+4$ for $u$ and $v$.